OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

will fan and medal!!

OpenStudy (anonymous):

|x - 2| + 10 =12 - 10 -10 -------------- |x - 2| = 2 x - 2 = 2 + 2 = +2 ------------ x = 4 is one of the solutions

OpenStudy (anonymous):

|x - 2 | + 10 = 12 x - 2 + 10 = 12 x - 8 = 12 +8 =+8 --------- x = 20 But I am pretty sure it should be positive not negative....hmm

OpenStudy (anonymous):

@yayanieves

OpenStudy (anonymous):

what do u think it is, can you find my mistake

OpenStudy (anonymous):

Idk. I really have no idea.@Chocolate1031

OpenStudy (anonymous):

But I still think it's C did u get it right? @yayanieves

OpenStudy (anonymous):

yeah I got it right:) @Chocolate1031

OpenStudy (anonymous):

thank goodness

OpenStudy (anonymous):

Lol thanks : )

OpenStudy (anonymous):

np ;) anytime